Two particles A and B with equal charges accelerat
Answer #1

Let the charge be q
Potential Energy Gained by particle A , = q*V
Potential Energy Gained by particle B , = q*4V

Using Energy Conservation,
1/2 * mA * vA^2 =  q*V ----1
1/2 * mB * vB^2 =  q*4V -----2

Dividing 1/2
mA * vA^2 /mB * vB^2 = 1/4
(4*mA)/mB = vB^2/vA^2

In Uniform Magnetic Field,
mA*vA^2/R = q*vA*B -----3

mB*vB^2/7R = q*vB*B -----4

Dividing 3/4
7*mA*vA/mB*vB = 1
(7*mA)/mB = vB/vA

Now using the two eq,

(4*mA)/mB = vB^2/vA^2
(4*mA)/mB = (7*mA)^2/mB^2
4 = 49*mA / mB

mA/mB = 4/49